OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of (f − g)(−4)? 4 3 −3 2

OpenStudy (anonymous):

(f-g)(x) = (x^2-13)-(x+5) so (f-g)(-4) = (-4)^2-13-(-4+5) = 16-13-1 =2

OpenStudy (anonymous):

THANK YOU SO MUCH!!!